### v1

# Contributions to conjectures in planar graphs: Induced Substructures, Treewidth, and Dominating Sets

Abstract
Two of the most prominent unresolved conjectures in graph theory, the Albertson-Berman conjecture and the Matheson-Tarjan conjecture, have been extensively studied by many researchers.
  • (AB)
    Every planar graph of order $n$ has an induced forest of order at least $\frac{n}{2}$.
  • (MT)
    Every plane triangulation of sufficiently large order $n$ has a dominating set of cardinality at most $\frac{n}{4}$.
Although partial results and weaker bounds than those originally conjectured have been obtained, both problems remain open. To contribute to their resolution, various generalizations and variations of the original concepts have been investigated, such as total dominating set, induced linear forests, and others.
In this paper, we clarify the relations among several notions related to these two major conjectures, such as connected domination and induced outerplanar subgraphs, etc., and survey the associated conjectures. We then provide counterexamples to some of these conjectures and establish the best bounds on the gap between the maximum orders of induced subgraphs under different structural conditions. In addition, we present a general upper bound on the order of induced subgraphs in terms of treewidth, a fundamental graph invariant.
Keywords and phrases. Induced subgraph, Treewidth, Connected domination, Planar graph, Outerplanar graph, $K_{4}$-Minor free graph, Triangulation, Albertson-Berman conjecture.
AMS 2020 Mathematics Subject Classification. 05C10, 05C35, 05C69.
1 Introduction
In graph theory, one of the most natural and fundamental questions is: Given a graph, how large can an induced particular subgraph be taken? Not only does this question represent a basic structural concern, but it also has implications in various contexts; for example, it may be applicable to the study of forbidden induced subgraphs, such as determining how large an induced forest must be forbidden in order to guarantee a desired property.
In this section, we discuss known results and conjectures relating to the above question divided into several subsections, while we introduce our results relevant to them. Then we summarize the relation among those unresolved conjectures. For the relevant definitions and terminology used throughout this paper, the reader is referred to Section 2.
1.1 Large induced forests in planar graphs
The Albertson-Berman conjecture [2] concerning large induced forests, which itself is of interest, remains unsolved. The maximun order of an induced forest in a graph $G$, denoted by $f(G)$, is called the induced forest number of $G$.
Conjecture A (Albertson and Berman [2])
Let $G$ be a planar graph of order $n$. Then $f(G)\geq\frac{n}{2}$.
If true, this would imply the existence of an independent set of cardinality at least $\frac{n}{4}$ in planar graphs, without relying on the Four-Color Theorem. Several partial results have been obtained as follows.
Theorem B (Hosono [19])
Let $G$ be an outerplanar graph of order $n$. Then $f(G)\geq\frac{2n}{3}$.
Theorem C (Le [21])
Let $G$ be a triangle-free planar graph of order $n$. Then $f(G)\geq\frac{5n}{9}$.
Generally, the best result concerning this conjecture is derived from Borodin’s work on the acyclic 5-colorability of planar graphs as follows.
Theorem D (Borodin [5])
Let $G$ be a planar graph of order $n$. Then $f(G)\geq\frac{2n}{5}$.
A related conjecture for bipartite planar graphs, posed by Akiyama and Watanabe [1], proposes a bound of $\frac{5n}{8}$, which is also unresolved.
Another variation by Chappell considers induced linear forests (which are forests consisting only of paths) in planar graphs, conjecturing a bound of $\frac{4n}{9}$ [25]. Similarly to Conjecture A, if true, this would also imply the existence of an independent set of cardinality at least $\frac{2n}{9}$. Recent progress has been made in this line of research. Notably, Dross, Montassier and Pinlou [13] proved the following result:
Theorem E (Dross et al. [13])
Let $G$ be a triangle-free planar graph of order $n$ and of size $m$. Then $G$ has an induced linear forest of order at least $\frac{9n-2m}{11}$.
Other recent developments include results on large induced subgraphs with bounded degree in outerplanar and planar graphs by D’Elia and Frati [12]. In particular, they proved that every planar graph of order $n$ has an induced subgraph with maximum degree at most 3 of order at least $\frac{5n}{13}$, and also constructed a planar graph of order $n$ whose largest induced subgraph with maximum degree at most 3 of order $\frac{4n}{7}+O(1)$.
1.2 Large induced forests in graphs with bounded treewidth
Another key structural parameter that measures how close a graph is to a tree or a forest is treewidth. A tree decomposition of a graph $G$ is a tree $T$ with vertices $X_{1},\dots,X_{n}$, where $X_{i}$ for each $i\in\{1,\dots,n\}$ is a subset of $V(G)$ such that
  1. (i)
    $\bigcup^{n}_{i=1}X_{i}=V(G)$,
  2. (ii)
    for each edge of $G$, there is an index $i$ such that both end vertices of the edge belong to $X_{i}$, and
  3. (iii)
    if both $X_{i}$ and $X_{j}$ contain a vertex $v$, then all nodes $X_{k}$ of $T$ in the unique path between $X_{i}$ and $X_{j}$ contain $v$ as well.
The width of a tree decomposition is defined as $\max_{i}|X_{i}|-1$, and the treewidth of $G$, denoted by $tw(G)$, is the minimum width of any tree decomposition of $G$. Note that $G$ is a forest if and only if $tw(G)=1$.
Intuitively, graphs with small treewidth tend to admit large induced forests. Indeed, Chappell and Pelsmajer [9] gave a lower bound on the order of induced forests in graphs with bounded treewidth.
Theorem F (Chappell and Pelsmajer [9])
Let $k\geq 2$ and $d\geq 2$ be integers, and $G$ be a graph of order $n$ with treewidth at most $k$. Then the maximum order of an induced forest in $G$ having maximum degree at most $d$ is at least $\frac{2dn+2}{kd+d+1}$ unless $G\in\{K_{1,1,3},K_{2,3}\}$ and $k=d=2$.
More generally, we can consider both the treewidth of a given graph and that of its induced subgraphs to derive the following result, which is the first result of ours.
Theorem 1
Let $s$ and $t$ be positive integers, and let $G$ be a graph with $tw(G)\leq s$. Then $G$ has an induced subgraph $H$ such that $tw(H)\leq t$ and $|H|\geq\frac{t+1}{s+1}|G|$.
In Section 3, we shall prove a theorem stronger than Theorem 1, and hence Theorem 1 follows as a corollary.
Theorem 1 has an important application as follows. If a planar graph $G$ has an induced subgraph of order at least $p$ that has treewidth at most 2, then $G$ has an induced forest of order $\frac{2p}{3}$ by Theorem 1. Thus, the existence of $K_{4}$-minor free induced subgraph of order at least $\frac{2n}{3}$ (Conjecture 3 described in the next subsection) improves the best known lower bound of Theorem D to $\frac{4n}{9}$.
Very recently, Matsumoto and Yashima [23] have investigated large induced subgraphs with given pathwidth, which is a path-variant of treewidth, in outerplanar graphs. This research is motivated by Chappell’s conjecture because every linear forest has pathwidth 1.
1.3 Large induced subgraphs concerning outerplanarity
From a broader perspective, although the Albertson-Berman and the Chappell conjectures (if true) would provide sharp lower bounds (with examples given in [1] and [25], respectively), one might expect to obtain larger lower bounds by considering a broader class of graphs, depending on the induced subgraphs under consideration. Indeed, Theorem 1 serves as an example of this approach. Thus, by focusing on the outerplanarity of graphs, we can consider the following invariants: the maximum order of an induced outerplanar (resp., outerplane, $K_{4}$-minor free) subgraph in a graph $G$, denoted by $s_{o}(G)$ (resp., $s_{o^{\prime}}(G)$, $s_{K_{4}}(G)$), is called the outerplanar (resp. outerplane, $K_{4}$-minor free) subgraph number of $G$. Among the above three invariants, the following relation holds.
Observation 2
For any plane graph $G$, $s_{o^{\prime}}(G)\leq s_{o}(G)\leq s_{K_{4}}(G)$.
In 2017, Borradaile, Le and Sherman-Bennett [4] showed that for every 2-outerplanar graph $G$, which is a planar graph close to an outerplanar graph, $s_{o}(G)\geq\frac{2n}{3}$. This together with Theorem B implies that such graphs admit an induced forest of order at least $\frac{4n}{9}$. Based on this, they proposed the following conjecture.
Conjecture G (Borradaile et al. [4])
Let $G$ be a planar graph of order $n$. Then $s_{o}(G)\geq\frac{2n}{3}$.
Motivated by their conjecture and Observation 2, we pose the following related conjecture.
Conjecture 3
Let $G$ be a planar graph of order $n$. Then $s_{K_{4}}(G)\geq\frac{2n}{3}$.
One may consider a similar problem for $s_{o^{\prime}}(G)$: for any plane graph $G$ of order $n$, does $s_{o^{\prime}}(G)\geq\frac{2n}{3}$ follow? However, this is not true in general. We construct a plane triangulation $G$ of order $n$ with $s_{o^{\prime}}(G)=\frac{5n}{8}+1$ in Section 8 (Remark 19). Generally, the following lower bound is known.
Theorem 4 (Angelini et al. [3])
Let $G$ be a plane graph of order $n$. Then $s_{o^{\prime}}(G)\geq\frac{n}{2}$.
Although the above two conjectures remain open, we establish the differences between $s_{o}(G)$ and $s_{K_{4}}(G)$, as well as between $s_{o^{\prime}}(G)$ and $s_{o}(G)$ as follows, which are the second results of ours and shall be proved in Sections 4 and 5, respectively.
Theorem 5
Let $G$ be a plane triangulation. Then the difference $s_{K_{4}}(G)-s_{o}(G)$ is arbitrarily large according to the order of $G$.
Theorem 6
Let $G$ be a plane triangulation. Then the difference $s_{o}(G)-s_{o^{\prime}}(G)$ is arbitrarily large according to the order of $G$.
Conversely, we also have two sharp upper bounds of those differences, which is the third result of ours.
Theorem 7
Let $G$ be a planar graph of order $n$. Then
$s_{K_{4}}(G)-s_{o}(G)\leq\frac{1}{5}s_{K_{4}}(G)\leq\frac{n}{5}.$
Furthermore, this bound is sharp.
Theorem 8
Let $G$ be a planar graph of order $n$. Then
$s_{o}(G)-s_{o^{\prime}}(G)\leq\frac{s_{o}(G)-2}{3}\leq\frac{n-2}{3}.$
Furthermore, this bound is sharp.
In Sections 6 and 7, we show that every $K_{4}$-minor free graph contains a large induced outerplanar or an outerplane subgraph, and then from those corresponding claims the above two theorems follow.
1.4 Dominations
As described above, the existence of large induced forests asserts the existence of a large maximal independent set in planar graphs. It is easy to see that every maximal independent set is a dominating set in a graph under consideration. In this subsection, we discuss relations among conjectures on large induced subgraphs and some on small vertex subsets such as a dominating set.
A subset $X$ of $V(G)$ is a dominating set of $G$ if $V(G)=\bigcup_{x\in X}N_{G}[x]$. The minimum cardinality of a dominating set of $G$, denoted by $\gamma(G)$, is called the domination number of $G$. For planar graphs, the Matheson-Tarjan conjecture [22] is another major open problem. They proved a partial result and provided examples showing that the bound $\frac{n}{4}$ would be tight if it were true.
Theorem H (Matheson and Tarjan [22])
Let $G$ be a plane triangulation of order $n$. Then $\gamma(G)\leq\frac{n}{3}$.
Conjecture I (Matheson and Tarjan [22])
Let $G$ be a plane triangulation of sufficiently large order $n$. Then $\gamma(G)\leq\frac{n}{4}$.
As a series of partial results on this, the following are known.
Theorem J (King and Pelsmajoer [20])
Let $G$ be a plane triangulation with maximum degree $6$. Then $\gamma(G)\leq\frac{n}{4}$.
Theorem K (Špacapan [27])
Let $G$ be a plane triangulation of order at least $7$. Then $\gamma(G)\leq\frac{17n}{53}$.
Theorem L (Christiansen et al. [10])
Let $G$ be a plane triangulation of order at least $11$. Then $\gamma(G)\leq\frac{2n}{7}$.
As with problems concerning induced subgraphs, domination has many variants. In our work, we focus on total domination and connected domination. A subset $X$ of $V(G)$ is a total dominating set of $G$ if $V(G)=\bigcup_{x\in X}N_{G}(x)$. The minimum cardinality of a total dominating set of $G$, denoted by $\gamma_{t}(G)$, is called the total domination number of $G$. A subset $X$ of $V(G)$ is a connected dominating set of $G$ if $V(G)=\bigcup_{x\in X}N_{G}[x]$ and the subgraph $G[X]$ is connected. The minimum cardinality of a connected dominating set of $G$, denoted by $\gamma_{c}(G)$, is called the connected domination number of $G$. Note that any connected dominating set consisting of at least two vertices is a total dominating set (for the study on domination invariants, see [17] for a recent survey, and [14, 15, 18] for books).
Similarly to Conjecture I, conjectures on these concepts for planar graphs exist.
Conjecture M (Bryant and Pavelescu [8])
Let $G$ be a plane triangulation of sufficiently large order $n$. Then $\gamma_{c}(G)\leq\frac{n}{3}$.
Conjecture N (Claverol et al. [11])
Let $G$ be a plane triangulation of order $n\geq 6$. Then $\gamma_{t}(G)\leq\frac{n}{3}$.
Interestingly, Conjecture M is equivalent to the following concerning the existence of a spanning tree with the maximum number of leaves, which is known as the maximum leaf number, due to Bradshaw, Masařík, Novotná and Stacho [7].
Conjecture O (Bradshaw et al. [7])
Let $G$ be a plane triangulation of sufficiently large order $n$. Then $G$ has a spanning tree with the maximum number of leaves at least $\frac{2n}{3}$.
In connection with the above, Noguchi and Zamfirescu [24] recently proved that any 4-connected plane triangulation of order $n$ has a spanning tree with exactly $k$ leaves for any integer $12\leq k\leq\frac{n}{2}+$ 1. Furthermore, they constructed a 4-connected plane triangulation of order $n$ which has no spanning tree with more than $\frac{2n}{3}$ leaves.
1.5 Relations among the conjectures and a counterexample
It is a notable observation that Conjecture M implies that $s_{o^{\prime}}\geq\frac{2}{3}|V(G)|$ for a plane triangulation $G$, and then it also implies Conjecture G: Take a plane triangulation $G$ and a connected dominating set $S$ (if necessary, we redraw $G$ on the plane so that a vertex in $S$ appears on the infinite face). Since $S$ is connected, it must be contained within a single face of $G-S$. Because $S$ dominates $G$, every vertex in $G-S$ must be adjacent to a vertex in $S$, and hence must appear on the boundary of that face. It follows that $G-S$ is outerplane.
The relations among the conjectures can be summarized in the following diagram (see Figure 1). In the diagram, $G$ denotes a plane triangulation of sufficiently large order $n$. Observe that Conjecture 3 implies Theorem H: Let $F$ be a largest induced $K_{4}$-minor free subgraph of $G$ and then $|F|\geq\frac{2n}{3}$. If $V(G)\setminus V(F)$ is not a dominating set, then there is a vertex $x\in V(F)$ such that $N_{G}[x]\subseteq V(F)$. However, since $G$ is a plane triangulation, $F[N_{G}[x]]$ trivially contains a $K_{4}$-minor, a contradiction. Thus, $V(G)\setminus V(F)$ is a dominating set.
[…]
Figure 1: Relations among the conjectures.
As can be seen from the diagram, Conjecture M is a crucial problem whose resolution would lead to the resolution of many other conjectures. However, we show that Conjecture M (i.e., Conjecture O) is false by providing an infinite family of counterexamples, which is the final result of ours and shall be proved in Section 8. In particular, by our construction, Theorem 9 follows even if the plane triangulation is Eulerian.
Theorem 9
There exists a family of plane triangulations $G$ such that $\gamma_{c}(G)\geq\frac{3|V(G)|}{8}-1$.
Therefore, Conjecture G cannot be proved via Conjecture M. However, the structural link between connected domination and induced outerplanar subgraphs remains valid. This leads us to propose a modified version of Conjecture M, reflecting the insights gained from our counterexamples.
Conjecture 10
Let $G$ be a plane triangulation of sufficiently large order $n$. Then $\gamma_{c}(G)\leq\frac{3n}{8}$.
Concerning this conjecture, the best known and partial results are the following.
Theorem P (Bose, Dujmović, Houdrouge, Morin and Odak [6])
Let $G$ be a plane triangulation of order $n\geq 3$. Then $\gamma_{c}(G)\leq\frac{10n}{21}$.
Theorem Q (Zhuang [28])
Let $G$ be a $K_{4}$-minor free graph of order $n\geq 4$, and $k$ be the number of vertices of degree $2$ in $G$. Then $\gamma_{c}(G)=1$ when $n\in\{4,5\}$; $\gamma_{c}(G)\leq\frac{2(n-k)}{3}$ when $n\geq 6$.
Although Conjecture M is not true, Conjectures G and 3 remain open. Thus, resolving them may lead to an improved lower bound for $f(G)$, and bring us closer to a complete resolution of Conjecture A.
2 Definitions and foundations
All graphs considered in this paper are finite, simple, and undirected. Let $G$ be a graph. We let $V(G)$ and $E(G)$ denote the set of vertices and the set of edges of $G$, respectively. For $x\in V(G)$, let $N_{G}(x)$, $N_{G}[x]$ and $d_{G}(x)$ denote the open neighborhood, the closed neighborhood, and the degree of $x$, respectively; thus $N_{G}(x)=\{y\in V(G)\mid xy\in E(G)\}$, $N_{G}[x]=N_{G}(x)\cup\{x\}$ and $d_{G}(x)=|N_{G}(x)|$. For a subset $X$ of $V(G)$, we let $G[X]$ denote the subgraph of $G$ induced by $X$. We let $K_{n}$ denote the complete graph of order $n$ and let $K_{m,n}$ denote the complete bipartite graph consisting of two partite sets of order $m$ and $n$. A vertex of degree $k$ is said to be a $k$-vertex, and a cycle of length $k$ is said to be a $k$-cycle.
A planar graph is a graph that can be drawn on the plane without edges crossing. If such a graph is already drawn on the plane, then the graph is called a plane graph. A plane triangulation $G$ is an edge-maximal plane graph, that is, for any non-adjacent vertices $u,v$, $G+uv$ is non-planar. An outerplanar graph is a planar graph that can be drawn on the plane such that all vertices appear on the boundary of the infinite face. Similarly to the plane graph, such a graph is called an outerplane graph if it is already drawn on the plane as above. Moreover, an outerplanar (or outerplane) graph is maximal if it is edge-maximal as well as a plane triangulation.
A minor of a graph $G$ is a graph that can be obtained from $G$ by removing vertices and removing or contracting edges. For a graph $H$, a graph $G$ is $H$-minor free if no minor of $G$ is isomorphic to $H$. A $K_{n}$-minor free graph $G$ is maximal if $G+uv$ is not a $K_{n}$-minor free graph for any two non-adjacent vertices $u$ and $v$ of $G$. For an integer $k\geq 1$, a $k$-tree is recursively defined as follows: A complete graph $K_{k+1}$ is a $k$-tree. Any graph obtained from a $k$-tree by adding a new vertex $v$ and making it adjacent to exactly $k$ vertices, such that they together $v$ induce $K_{k+1}$, is also a $k$-tree. It is well known that the maximal $K_{4}$-minor free graphs are exactly the 2-trees. Note that a maximal outerplanar graph must be a maximal $K_{4}$-minor free graph, i.e., a 2-tree.
3 Proof of Theorem 1
In this section, we prove the following theorem stronger than Theorem 1.
Theorem 11
Every $k$-tree $G$ has the unique partition of $V(G)$ into $k+1$ independent sets for any integer $k\geq 1$. Moreover, for any integer $12\leq t\leq k+$ 1, the union of any $t$ subsets in this partition induces a $(t-1)$-tree.
Proof.
Let $G$ be a $k$-tree for an integer $k\geq 1$. Let $K_{k+1}=G_{0},G_{1},\dots,G_{m}=G$ be a sequence of $k$-trees such that for each $10\leq i\leq m-$ 1, $G_{i+1}$ is obtained from $G_{i}$ by adding a new vertex and joining it to exactly $k$ vertices of a $K_{k}$ in $G_{i}$. We first show by induction on $m$ that $G$ has a unique partition of $V(G)$ into $k+1$ independent sets. When $m=0$, the claim trivially holds. So suppose $m\geq 1$. Let $v\in V(G)\setminus V(G_{m-1})$ and $S$ be the set of $k$ vertices adjacent to $v$ in $G$. By the induction hypothesis, $G_{m-1}$ has the unique partition $U^{\prime}_{1},U^{\prime}_{2},\dots,U^{\prime}_{k+1}$ of $V(G_{m-1})$ each of which is independent. Without loss of generality, we may suppose that $|S\cap U^{\prime}_{i}|=1$ for each $1\leq i\leq k$ and $S\cap U^{\prime}_{k+1}=\emptyset$. Let $U_{1},U_{2},\dots,U_{k+1}$ be a partition of $V(G)$ each of which is independent. Similarly to the above, we may suppose that $S\cap U_{i}=S\cap U^{\prime}_{i}$ for each integer $1\leq i\leq k$ and $S\cap U_{k+1}=\emptyset$. Since $V(G_{m-1})$ is uniquely partitionable, the partition $U_{1},U_{2},\dots,U_{k+1}$ is uniquely determined in $V(G)\setminus\{v\}$. Moreover, since $v$ is adjacent to all vertices in $S$, $v$ must be in $U_{k+1}$. Thus, the partition $U_{1},U_{2},\dots,U_{k+1}$ is uniquely determined.
Next, we show that for any integer $12\leq t\leq k+$ 1, the union of any $t$ subsets in the unique partition induces a $(t-1)$-tree $H$. Let $U_{1},U_{2},\dots,U_{k+1}$ be the unique partition of $V(G)$ each of which is independent. It suffices to prove that $G^{\prime}=G[V(G)\setminus U_{k+1}]$ is a $(k-1)$-tree. We consider the process of building $G$, say $K_{k+1}=G_{0},G_{1},\dots,G_{m}=G$. Let $K_{k}=G^{\prime}_{0},G^{\prime}_{1},\dots,G^{\prime}_{m}=G^{\prime}$ be the process of building $G^{\prime}$ obtained from the above process of $G$ by deleting vertices in $U_{k+1}$. If $G_{i+1}$ is obtained from $G_{i}$ by adding a vertex $v$ in $U_{j}$ for $j\in\{1,2,\dots,k\}$, then $v$ is adjacent to exactly $k-1$ vertices in $G^{\prime}_{i+1}$ in which they together $v$ form a $K_{k}$. Otherwise, if $G_{i+1}$ is obtained from $G_{i}$ by adding a vertex $v$ in $U_{k+1}$, we have $G^{\prime}_{i+1}=G^{\prime}_{i}$. Thus, every graph generated in the process $K_{k}=G^{\prime}_{0},G^{\prime}_{1},\dots,G^{\prime}_{m}=G^{\prime}$ is a $(k-1)$-tree. ∎
It is well known that a $k$-tree is an edge-maximal graph with treewidth $k$, i.e., every graph with treewidth $k$ is a subgraph of a $k$-tree. By Theorem 11, if a graph $G$ has treewidth at most $s$, $G$ has a partition of $V(G)$ into $s+1$ independent sets $U_{1},U_{2},\dots,U_{s+1}$ with $|U_{1}|\geq|U_{2}|\geq\cdots\geq|U_{s+1}|$ (where some of them are possibly empty). Every $t+1$ subsets in the partition induces a graph with treewidth at most $t$ by Theorem 11. Therefore, an induced subgraph $H=G[U_{1}\cup U_{2}\cup\cdots\cup U_{t+1}]$ has treewidth at most $t$ and its order is at least $\frac{(t+1)|V(G)|}{s+1}$, and then Theorem 1 follows.
4 Proof of Theorem 5
We show that the difference $s_{K_{4}}(G)-s_{o}(G)$ is arbitrarily large according to the order of a plane triangulation $G$. To see this, we construct a plane triangulation $D_{k}$ such that $s_{K_{4}}(D_{k})\geq\frac{12(|V(D_{k})|-1)}{20}$ and $s_{o}(D_{k})\leq\frac{23(|V(D_{k})|-1)}{40}$.
Construction: We first prepare a plane embedding $O$ of the octahedron graph and insert a single vertex $v$ into a face of $O$ joined to all the vertices on the boundary of the face. The resulting graph is denoted by $R^{\prime}$ and the boundary of the face consisting of all vertices of degree $4$ is denoted by $abc$ (see Figure 2). Next we prepare a $K_{4}$ on the plane, where $V(K_{4})=\{x_{0},x_{1},x_{2},x_{3}\}$. For each integer $i\in\{1,2,3\}$, let $F_{i}$ denote the face bounded by $x_{i},x_{i+1}$ and $x_{0}$, where the indices are taken modulo $3$. We replace $F_{i}$ with a copy of $R^{\prime}$, identifying the boundary $3$-cycle of $F_{i}$ with the boundary cycle $abc$ of $R^{\prime}$. The resulting graph is denoted by $R$. Note that $R$ can be partitioned into exactly four $K_{4}$s.
$a$$b$$c$$v$
Figure 2: The graph $R^{\prime}$.
Let $k$ be an even integer with $k\geq 4$, and let $R_{1},\dots,R_{k}$ be disjoint copies of $R$. For each integer $1\leq j\leq k$, let $x_{j,0},x_{j,1},x_{j,2}$ and $x_{j,3}$ be the vertices of $R_{j}$ corresponding to $x_{0},x_{1},x_{2}$ and $x_{3}$ of $R$, respectively. We now define the graph $D_{k}$ (see Figure 3) by
[…]
$E(D_{k})$ =$=\{x_{i,1}x_{i+1,1},x_{i,3}x_{i+1,1},x_{i,3}x_{i+1,2}\mid 1\leq i\leq k-1\}$
$\quad\cup\{xz\,|\,x\in\bigcup_{1\leq j\leq k}\{x_{j,1},x_{j,2},x_{j,3}\}\}\cup\bigcup_{1\leq j\leq k}E(R_{j}).$
$z$$x_{1,2}$$x_{1,3}$$x_{1,1}$$x_{1,0}$$x_{2,2}$$x_{2,3}$$x_{2,1}$$x_{2,0}$$x_{k-1,2}$$x_{k-1,3}$$x_{k-1,1}$$x_{k-1,0}$$x_{k,2}$$x_{k,3}$$x_{k,1}$$x_{k,0}$
Figure 3: The graph $D_{k}$, where a copy of $R^{\prime}$ is embedded into each shaded triangle.
We first consider $s_{o}(D_{k})$. Let $H$ be a largest induced outerplanar subgraph of $D_{k}$. Since every outerplanar graph is $K_{4}$-minor free, $H$ takes at most three vertices of each $K_{4}$ subgraph of $D_{k}$, that is, $|V(H)\cap V(R_{j})|\leq 3\cdot 4=12$ for each integer $1\leq j\leq k$. Moreover, if $|V(H)\cap V(R_{j})|=12$, then $x_{j,1},x_{j,2},x_{j,3}\in V(H)\cap V(R_{j})$; otherwise, for example, if $x_{j,1},x_{j,2},x_{j,0}\in V(H)\cap V(R_{j})$, then they together with three vertices in the triangular region $x_{j,1}x_{j,2}x_{j,0}$ belonging to $V(H)\cap V(R_{j})$ contains a $K_{4}$-minor. Moreover, if there exists $11\leq i\leq k-$ 1 such that $|V(H)\cap V(R_{i})|=12$ and $|V(H)\cap V(R_{i+1})|=12$, $\{x_{i,1},x_{i,2},x_{i,3},x_{i+1,1},x_{i+1,2},x_{i+1,3}\}$ and three vertices of $H$ in the interior of the region bounded by $x_{i,0}x_{i,1}x_{i,3}$ induce a $K_{2,3}$-minor. Thus, for each odd integer $11\leq j\leq k-$ 1, $|V(H)\cap V(R_{j})|+|V(H)\cap V(R_{j+1})|\leq 12+11=23$, that is, $s_{o}(D_{k})=|V(H)|\leq 23\cdot\frac{k}{2}=\frac{23k}{2}$. (Note that $z\notin V(H)$ if there is an index $j$ such that $|V(H)\cap V(R_{j})|=12$.)
Next, we consider $s_{K_{4}}(D_{k})$. We see that $M=D_{k}[\bigcup_{1\leq j\leq k}\{x_{j,1},x_{j,2},x_{j,3}\}]$ is (maximal) outerplanar. Moreover, for each $R^{\prime}$ placed in a triangular region $x_{j,0}x_{j,i}x_{j,i^{\prime}}$ ($1\leq j\leq k$, $i,i^{\prime}\in\{1,2,3\}$, $i\neq i^{\prime}$), the central vertex $v_{c}$ corresponding to $v$ and other two vertices adjacent to $v_{c}$ together with $x_{j,1},x_{j,2},x_{j,3}$ do not contain any $K_{4}$-minor (since the induced subgraph is obtained from a triangle $x_{j,1}x_{j,2}x_{j,3}$ by adding a vertex of degree 2 repeatedly). Let $W_{j}$ be the set of such nine vertices in total, and let $W=\bigcup_{1\leq i\leq k}W_{i}$. Since the whole induced subgraph $Q=D_{k}[V(M)\cup W]$ is a $K_{4}$-minor free graph, we have $s_{K_{4}}(D_{k})\geq|V(Q)|=12k$. Therefore, the difference $s_{K_{4}}(D_{k})-s_{o}(D_{k})$ is arbitrarily large according to the order of the triangulation.
5 Proof of Theorem 6
We show that the difference $s_{o}(G)-s_{o^{\prime}}(G)$ is arbitrarily large according to the order of a plane triangulation $G$. To see this, we construct a plane triangulation $G_{k}$ in which there exists an outerplanar subgraph $H$ such that $s_{o}(G_{k})\geq\frac{15}{22}|V(G_{k})|$, and an outerplane subgraph $H^{\prime}$ such that $s_{o^{\prime}}(G_{k})\leq\frac{7}{11}|V(G_{k})|$.
Construction: For any integer $k\geq 4$, we construct the graph $G_{k}$ as follows. We first prepare a plane embedding $O$ of the octahedron graph and replace an inner face $F$ of $O$ with a copy of another octahedron $O^{\prime}$ by identifying the $3$-cycle of $F$ with the boundary cycle of $O^{\prime}$. The resulting graph is denoted by $I$. Moreover, set $S=\{v_{1},w_{1},w_{2},w_{3}\}$ (see Figure 4).
$u_{1}$$u_{2}$$u_{3}$$v_{1}$$v_{2}$$v_{3}$$w_{1}$$w_{2}$$w_{3}$
Figure 4: The graph $I$, where the four vertices enclosed in circles form the set $S$.
We next prepare a $K_{4}$ with vertex set $V(K_{4})=\{x_{0},x_{1},x_{2},x_{3}\}$, embedded on the plane so that the outer face is bounded by the $3$-cycle $x_{1}x_{2}x_{3}$. For each integer $1\leq i\leq 3$, let $F_{i}$ denote the face bounded by $x_{i},x_{i+1}$ and $x_{0}$, where the indices are taken modulo $3$. We replace $F_{i}$ with a copy of $I$, identifying the boundary $3$-cycle of $F_{i}$ with the boundary cycle $u_{1}u_{2}u_{3}$ of $I$. The resulting graph is denoted by $H$.
Let $k$ be an integer with $k\geq 4$, and let $H_{1},\dots,H_{k}$ be disjoint copies of $H$. For each integer $1\leq j\leq k$, let $x_{j,1},x_{j,2},x_{j,3}$ and $x_{j,0}$ be the vertices of $V(H_{j})$ which correspond to $x_{1},x_{2},x_{3}$ and $x_{0}$, respectively. We now define the graph $G_{k}$ (see Figure 5) by
[…]
$E(G_{k})$ =$=\{x_{i,1}x_{i+1,1},x_{i,3}x_{i+1,1},x_{i,3}x_{i+1,2},x_{i,1}x_{k,3}\mid 1\leq i\leq k-1\}$
$\hskip 14.22636pt\cup\{x_{2,3}x_{i,2},x_{2,3}x_{i,3}\mid 3\leq i\leq k\}\cup\{x_{1,2}x_{k,3},x_{1,2}x_{2,3},x_{1,3}x_{2,3}\}\cup\bigcup_{1\leq j\leq k}E(H_{j}).$
$x_{1,2}$$x_{1,3}$$x_{1,1}$$x_{1,0}$$x_{2,2}$$x_{2,3}$$x_{2,1}$$x_{2,0}$$x_{k-1,2}$$x_{k-1,3}$$x_{k-1,1}$$x_{k-1,0}$$x_{k,2}$$x_{k,3}$$x_{k,1}$$x_{k,0}$
Figure 5: The graph $G_{k}$, where a copy of $I$ is embedded into each shaded triangle.
We first consider $s_{o^{\prime}}(G_{k})$. Let $Q$ be a largest induced outerplane subgraph of $G_{k}$. Similarly to the proof of Theorem 5, $Q$ takes at most four vertices from an octahedron graph of $I$. Moreover, $Q$ takes at most two vertices can be selected from each $K_{4}$ of $G_{k}$: otherwise, for example, if $Q$ takes three vertices $x_{j,1},x_{j,2},x_{j,3}$, then $Q$ cannot take any vertex from the octahedron graph in the interior of the region bounded by $x_{j,1}x_{j,2}x_{j,3}$, which contradicts to the maximality of $Q$. Therefore, we have $s_{o^{\prime}}(G_{k})\leq(4\cdot 3+2)k=14k=\frac{7}{11}|V(G_{k})|$; note that $|V(G_{k})|=22k$.
Next, we consider $s_{o}(G_{k})$. For each integer $1\leq i\leq 3$ and each integer $1\leq j\leq k$, let $S_{j,i}$ denote the subset of $V(H_{j})$ which corresponds to $S$ in $F_{j,i}$, where $F_{j,i}$ denotes a region of $H_{j}$ correponding to $F_{i}$, and set $X_{j}=\{x_{j,1},x_{j,2},x_{j,0}\}$. Now the induced subgraph
$R=G_{k}\left[\left(\bigcup_{1\leq j\leq k,1\leq i\leq 3}S_{j,i}\right)\cup\left(\bigcup_{1\leq j\leq k}X_{j}\right)\right]$
is clearly outerplanar; note that for any pair of $i,j$ with $i\neq j$, there is only one edge $x_{i,1}x_{j,1}$ between $H_{i}$ and $H_{j}$. Therefore, we have $s_{o}(G_{k})\geq(4\cdot 3+3)k=15k=\frac{15}{22}|V(G_{k})|$.
Hence, the graph $G_{k}$ shows that the difference $s_{o}(G_{k})-s_{o^{\prime}}(G_{k})$ is arbitrarily large according to the order.
6 Proof of Theorem 7
To prove Theorem 7, we give a sharp lower bound on $s_{o}(G)$ in terms of $s_{K_{4}}(G)$.
Theorem 12
Let $G$ be a planar graph. Then, $\frac{4}{5}s_{K_{4}}(G)\leq s_{o}(G)$.
Note that we can immediately obtain Theorem 7 by Theorem 12. There exist graphs that attain this bound, which are introduced at the end of this section.
To prove Theorem 12, it suffices to show the following lemma.
Lemma 13
Any $K_{4}$-minor free graph $G$ of order $n$ has an induced outerplanar subgraph of order at least $\frac{4n}{5}$.
Before proving Lemma 13, we prepare some definitions and notations. For a positive integer $k$, a $k$-clique in a graph $G$ is a clique of order $k$ in $G$. For a $2$-tree $G$, we define a graph $\Delta_{G}$ as follows: The vertex set of $\Delta_{G}$ consists of the $3$-cliques in $G$. Two vertices of $\Delta_{G}$ are adjacent if they share a common pair of adjacent vertices of $G$. Note that every spanning tree of $\Delta_{G}$ can be regarded as a tree decomposition of $G$. If $G$ is (maximal) outerplanar, then $\Delta_{G}$ is a tree with maximum degree at most three. Moreover, if $G$ is already embedded on the plane as an outerplane graph, then $\Delta_{G}$ corresponds to the graph obtained from the dual of $G$ by removing the vertex corresponding to the outerface of $G$.
The following lemma clearly holds.
Lemma 14
If every block of a graph $G$ is outerplanar, then $G$ is also outerplanar.
Outerplanar graphs can be characterized as graphs that contain neither $K_{4}$ nor $K_{2,3}$ as minors, while for $2$-trees, outerplanarity can be determined by whether $K_{1,1,3}$ is contained as a subgraph or not, where $K_{1,1,3}$ denotes the graph obtained from $K_{2}$ by adding three independent vertices and joining them to all vertices of the $K_{2}$.
Lemma 15
A $2$-tree $G$ is not maximal outerplanar if and only if $G$ has a subgraph isomorphic to $K_{1,1,3}$.
Proof.
of Lemma 15 It is clear that if a graph $G$ has $K_{1,1,3}$ as a subgraph, then $G$ is not outerplanar. Hence, we shall show the necessity. We suppose that a $2$-tree $G$ has no $K_{1,1,3}$ as a subgraph. Then, each edge of $G$ is contained in at most two triangles. We shall show that $G$ is a maximal outerplanar graph by induction on the order of $G$. Since every $2$-tree of order at most $4$ is outerplanar, we may assume that $G$ has at least $5$ vertices. Then, there exists a vertex $a$ of degree $2$ in $G$. Let $N_{G}(a)=\{b,c\}$. Let $G^{\prime}$ be the graph obtained from $G$ by removing $a$. Then $G^{\prime}$ is also a $2$-tree and the edge $bc$ belongs to at most one triangle in $G^{\prime}$. By induction hypothesis, $G^{\prime}$ can be embedded on the plane so that all vertices of $G^{\prime}$ appear on the boundary of the outerface. In this embedding, the edge $bc$ belongs to the boundary of the outerface. In this setting, adding the vertex $a$ on the outerface of $G^{\prime}$ and adding edges $ab$ and $ac$, we can obtain an embedding of $G$, which is an outerplane graph. ∎
For a rooted tree $(T,r)$, if a vertex $v$ in $(T,r)$ has no child, then we call $v$ a leaf, and if $v$ has at least one child and all its children are leaf, then we call $v$ a pre-leaf.
Proof.
of Lemma 13 Let $G$ be a counterexample to Lemma 13 with the smallest number of vertices. We may assume that $G$ is a $2$-tree. Non-outerplanar $2$-tree of order at most $5$ must be $K_{1,1,3}$, which has an induced outerplanar subgraph of order $4$. Thus, we may assume that the order of $G$ is at least $6$.
Claim 6.1
For any vertex $v$ of $G$, the neighborhood of $v$ contains at most three $2$-vertices.
Proof.
Suppose that $N_{G}(v)$ contains four $2$-vertices $u_{1},u_{2},u_{3}$ and $u_{4}$. Let $G^{\prime}=G-\{v,u_{1},u_{2},u_{3},u_{4}\}$. By the minimality of $G$, there exists an induced outerplanar subgraph $H^{\prime}$ of order at least $\frac{4}{5}|V(G^{\prime})|$ in $G^{\prime}$. The induced subgraph $H=G[V(H^{\prime})\cup\{u_{1},u_{2},u_{3},u_{4}\}]$ can be obtained from $H^{\prime}$ by adding pendant vertices $u_{1},u_{2},u_{3}$ and $u_{4}$. Hence, $H$ is outerplanar and we have
$|V(H)|=|V(H^{\prime})|+4\geq\frac{4}{5}|V(G^{\prime})|+4=\frac{4}{5}(|V(G^{\prime})|+5)=\frac{4}{5}|V(G)|,$
which is a contradiction. ∎
We choose a spanning tree $T$ of $\Delta_{G}$ and a root $R$ so that the number of leaves is maximized. Note that we have $|V(T)|\geq 4$ since $|V(G)|\geq 6$. Let $P=\{a,b,c\}$ be a pre-leaf of $(T,R)$.
Claim 6.2
Every vertex in $(T,R)$ has at most four leaves as children.
Proof.
If a vertex $Q$ in $(T,R)$ has at least five leaves as children, then at least one of the vertices in the $3$-clique $Q$ is adjacent to at least four $2$-vertices in $G$, which contradicts Claim 6.1. ∎
Claim 6.3
$P$ is not the root $R$.
Proof.
Suppose that $P$ is the root $R$. It follows from Claim 6.2 that $P$ has at most four leaves as children, and hence $|V(T)|\leq 5$, that is, $|V(G)|\leq 7$. As $|V(G)|\geq 6$, $G$ is isomorphic to $G_{1}$ or $G_{2}$ as shown in Figure 6. However, $G_{1}$ and $G_{2}$ have induced outerplanar subgraphs of order $5$ and $6$, respectively, a contradiction. ∎
Figure 6: Two graphs $G_{1}$ and $G_{2}$.
Claim 6.4
If a $3$-clique $P^{\prime}=\{v_{1},v_{2},v_{3}\}$ in $G$ is neither the root $R$ nor a leaf in $(T,R)$, then $d_{G}(v_{i})\geq 3$ for any $1\leq i\leq 3$.
Proof.
Let $P^{\prime}=\{v_{1},v_{2},v_{3}\}$ be a vertex in $(T,R)$ which is neither the root nor a leaf. Suppose that $d_{G}(v_{1})=2$. Then, the parent and every child of $P^{\prime}$ include $\{v_{2},v_{3}\}$, that is, the parent and every child of $P^{\prime}$ are adjacent in $\Delta_{G}$. For each children $C$ of $P^{\prime}$, removing the edge $P^{\prime}C$ and adding the edge between $C$ and the parent of $P^{\prime}$, we can obtain another spanning tree $T^{\prime}$ of $\Delta_{G}$. That has one more leaf than $T$, which contradicts the choice of $(T,R)$. ∎
Let $Q$ be the parent of the pre-leaf $P=\{a,b,c\}$. We may suppose that both $P$ and $Q$ include $\{a,b\}$, and let $d$ be the vertex in the $3$-clique $Q$ but not in $P$; Thus, $Q=\{a,b,d\}$.
Claim 6.5
$3\leq d_{G}(c)\leq 5$.
Proof.
Since $P$ is pre-leaf, it follows from Claim 6.4 that $d_{G}(c)\geq 3$. We suppose that $d_{G}(c)\geq 6$. Since every neighbor of $c$ other than $a$ and $b$ has degree $2$, $c$ is adjacent to at least four $2$-vertices, which contradicts Claim 6.1. ∎
Let $X$ be the set of $2$-vertices of $G$ whose neighborhood is included in $P$ and that is not $d$. Let $X$ consist of $k$ vertices $x_{1},x_{2},\ldots,x_{k}$. Then, $P$ has $k$ children. Since $P$ is pre-leaf, it follows from Claim 6.2 that $1\leq k\leq 4$. By Claim 6.5, we have $d_{G}(c)\geq 3$, and hence there is a vertex in $X$ adjacent to $c$. Thus, we may assume that $N_{G}(x_{1})=\{a,c\}$, that is, a $3$-clique $\{a,c,x_{1}\}$ is a child of $P$. In this setting, $G$ has the structure shown in Figure 7, where shaded areas represent subgraphs induced by the vertices that are not contained in $P\cup Q\cup X$. (Some of the vertices $x_{2},x_{3}$ and $x_{4}$ in Figure 7 may not exist, and their neighbors may be different.)
Figure 7: The structure of $G$.
Claim 6.6
There are exactly two two children of $P$, that is, $k=2$, and $N_{G}(x_{1})=N_{G}(x_{2})=\{a,c\}$.
Proof.
First, suppose that there are only two $3$-cliques including $\{a,c\}$: $P$ and $\{a,c,x_{1}\}$. Then, there is no subgraph having $x_{1}$ and isomorphic to $K_{1,1,3}$. Let $G^{\prime}=G-x_{1}$. By the minimality of $G$, there exists an induced outerplanar subgraph $H^{\prime}$ of order at least $\frac{4}{5}|V(G^{\prime})|$ in $G^{\prime}$. The induced subgraph $H=G[V(H^{\prime})\cup\{x_{1}\}]$ has no $K_{1,1,3}$ as a subgraph, and hence $H$ is outerplanar by Lemma 15, while $|V(H)|=|V(H^{\prime})|+1>\frac{4}{5}(|V(G^{\prime})|+1)=\frac{4}{5}|V(G)|$, a contradiction. Therefore, $k\geq 2$ and we may assume that $P$ has a child $\{a,c,x_{2}\}$, that is, $N_{G}(x_{2})=\{a,c\}$.
Second, we suppose that $k\geq 3$. It follows from Claim 6.1 that there is at most two $3$-cliques including $\{b,c\}$, one of which is $P$. Then, there is no subgraph having the edge $bc$ and isomorphic to $K_{1,1,3}$. Let $G^{\prime}=G-(X\cup\{a,c\})$. By the minimality of $G$, there exists an induced outerplanar subgraph $H^{\prime}$ of order at least $\frac{4}{5}|V(G^{\prime})|$ in $G^{\prime}$. The induced subgraph $H=G[V(H^{\prime})\cup X\cup\{c\}]$ has no $K_{1,1,3}$ as a subgraph and hence $H$ is outerplanar by Lemma 15. However, since $k\geq 3$, we have
$|V(H)|=|V(H^{\prime})|+k+1\geq\frac{4}{5}\left(|V(G^{\prime})|+k+\frac{1}{4}k+\frac{5}{4}\right)\geq\frac{4}{5}\left(|V(G^{\prime})|+k+2\right)\geq\frac{4}{5}|V(G)|,$
which is a contradiction. ∎
Since every pre-leaf in $(T,R)$ is not the root by Claim 6.3, there exists a vertex of $(T,R)$ which is the parent of a pre-leaf and whose every child is either a leaf or a pre-leaf. Hereafter, we suppose that $Q$ is such a vertex.
Claim 6.7
$Q$ has no leaf as a child.
Proof.
Suppose that $Q$ has a leaf as a child. Then, there exists a $2$-vertex $y$ of $G$ whose neighborhood is included in $Q$. Let $Y=\{c,x_{1},x_{2},y\}$, and if $N_{G}(y)$ contains $a$, we let $Z=Y\cup\{a\}$, and otherwise let $Z=Y\cup\{b\}$. Let $G^{\prime}=G-Z$. By the minimality of $G$, there exists an induced outerplanar subgraph $H^{\prime}$ of order at least $\frac{4}{5}|V(G^{\prime})|$ in $G^{\prime}$. It is easy to see that the induced subgraph $H=G[V(H^{\prime})\cup Y]$ has no $K_{1,1,3}$ as a subgraph and hence $H$ is outerplanar by Lemma 15. However, we have $|V(H)|=|V(H^{\prime})|+4\geq\frac{4}{5}|V(G)|$, which is a contradiction. ∎
By Claim 6.7, all children of $Q$ are pre-leaves.
Claim 6.8
$Q$ has at least two children.
Proof.
Suppose that $Q$ has only one child $P$. If $Q$ is the root $R$, then by Claim 6.8, $G$ is isomorphic to the graph as shown in Figure 8, which has an induced outerplanar subgraph of order $5$, a contradiction. Thus, we may assume that $Q$ is not the root $R$.
Figure 8: A $2$-tree of order $6$.
Let $Q^{\prime}$ be the parent of $Q$. It follows from Claim 6.4 that $Q^{\prime}$ contains $d$, since otherwise $d_{G}(d)$ would be $2$. Let $Z=\{a,b,c,x_{1},x_{2}\}$, and if $Q\cap Q^{\prime}=\{a,d\}$, we let $Y=\{b,c,x_{1},x_{2}\}$, and if $Q\cap Q^{\prime}=\{b,d\}$ let $Y=\{a,c,x_{1},x_{2}\}$. Let $G^{\prime}=G-Z$. By the minimality of $G$, there exists an induced outerplanar subgraph $H^{\prime}$ of order at least $\frac{4}{5}|V(G^{\prime})|$ in $G^{\prime}$. It is easy to see that the induced subgraph $H=G[V(H^{\prime})\cup Y]$ has no $K_{1,1,3}$ as a subgraph, and hence $H$ is outerplanar by Lemma 15. However, we have $|V(H)|=|V(H^{\prime})|+4\geq\frac{4}{5}|V(G)|$, which is a contradiction. ∎
Let $P^{\prime}$ be a child of $Q$ which is not $P$, and $c^{\prime}$ be the vertex in $P^{\prime}$ but not in $Q$. Similarly to $P$, we may assume that $P^{\prime}$ has two children. Then, there exists two $2$-vertices $x^{\prime}_{1}$ and $x^{\prime}_{2}$ such that $N_{G}(x^{\prime}_{1})=N_{G}(x^{\prime}_{2})\subset P^{\prime}$. It follows from Claim 6.1 that we only have to consider the following four situations: $(1)$ to $(4)$, as shown in Figure 9.
Figure 9: The four possible situations around $P$ and $P^{\prime}$.
Let $Y=\{c,c^{\prime},x_{1},x_{2},x^{\prime}_{1},x^{\prime}_{2}\}$, and if the case is type $(1)$, let $Z=Y\cup\{a\}$, and otherwise, let $Z=Y\cup\{b\}$. Let $G^{\prime}=G-Z$. By the minimality of $G$, there exists an induced outerplanar subgraph $H^{\prime}$ of order at least $\frac{4}{5}|V(G^{\prime})|$ in $G^{\prime}$. It is easy to see that the induced subgraph $H=G[V(H^{\prime})\cup Y]$ has no $K_{1,1,3}$ as a subgraph, and hence $H$ is outerplanar by Lemma 15. However, we have $|V(H)|=|V(H^{\prime})|+6>\frac{4}{5}(|V(G^{\prime})|+5)=\frac{4}{5}|V(G)|$, which is a contradiction. Therefore, the proof of Lemma 13 is finished. ∎
Now we construct a family of $K_{4}$-minor free graphs that shows that the bound given by Theorem 12 is tight. We take several disjoint copies of $K_{1,1,3}$ (or $K_{2,3}$) and add any number of edges between them, as long as the resulting graph remains $K_{4}$-minor-free. For example, a graph shown in Figure 10 is such a construction.
Figure 10: A $2$-tree attaining the bound given by Theorem 12.
7 Proof of Theorem 8
To prove Theorem 8, it is sufficient to prove the following.
Theorem 16
Let $G$ be a $K_{4}$-minor free graph of order $n$. For any embedding of $G$ on the plane, $G$ has an induced outerplane subgraph of order at least $\lceil\frac{2n+2}{3}\rceil$.
This bound can even be attained by outerplanar graphs, which is introduced after proving Theorem 16. It follows from Theorem 16 and its example that the following corollary provides the best possible bound, which leads to Theorem 8.
Corollary 17
Let $G$ be a plane graph of order $n$. Then,
$\left\lceil\frac{2s_{K_{4}}(G)+2}{3}\right\rceil\leq\left\lceil\frac{2s_{o}(G)+2}{3}\right\rceil\leq s_{o^{\prime}}(G).$
Let $H$ be a plane graph. A cycle $C$ in $H$ is separating if there exist vertices $p$ and $q$ of $H$ such that $p$ and $q$ lie inside $C$ and outside $C$, respectively. In this situation, we say that $C$ separates $p$ and $q$. Note that $H$ is an outerplane graph if and only if there is no separating cycle in $H$.
Proof.
of Theorem 16 It suffices to verify the theorem for $2$-trees. Let $G$ be a $2$-tree of order $n$ embedded on the plane. The case where $G$ has at most five vertices can be easily checked. Thus, we may assume that $G$ has at least six vertices.
By Theorem 11, $V(G)$ can be partitioned into three independent sets $V_{1},V_{2}$ and $V_{3}$ so that the union of any two of them induced a forest. We may assume that $|V_{1}|\geq|V_{2}|\geq|V_{3}|$. It is easy to see that the inequality $|V_{1}\cup V_{2}|>\lceil\frac{2n+2}{3}\rceil$ holds if $|V_{2}|>|V_{3}|$. Moreover, it also holds if $|V_{2}|=|V_{3}|$ and $|V_{1}|\geq|V_{2}|+2$. Since any embedding of a forest is an outerplane graph, the induced subgraph $G[V_{1}\cup V_{2}]$ is an outerplane graph. Thus, we may assume that $|V_{2}|=|V_{3}|$ and $|V_{2}|\leq|V_{1}|\leq|V_{2}|+1$. Let $V_{3}$ consist of $m$ vertices $x_{1},x_{2},\ldots,x_{m}$. Then, $|V_{2}|=m$ and $m\leq|V_{1}|\leq m+1$. Note that $m\geq 2$ since $G$ has at least six vertices. In this setting, we have $|V_{1}|+|V_{2}|+1\geq\lceil\frac{2n+2}{3}\rceil$. Thus, we may assume that there is no vertex $x$ in $V_{3}$ such that the induced subgraph $G[V_{1}\cup V_{2}\cup\{x\}]$ is an outerplane graph; otherwise, the proof is complete.
For each integer $1\leq i\leq m$, the induced subgraph $H_{i}=G[V_{1}\cup V_{2}\cup\{x_{i}\}]$ is not an outerplane graph, and hence there exists a separating cycle in $H_{i}$. Furthermore, we now show that such a cycle can be chosen to have length $3$. Suppose that there exists a separating cycle $\ell^{\prime}_{i}$ in $H_{i}$ whose length is at least $4$. The vertices of $\ell^{\prime}_{i}$ other than $x_{i}$ are contained in $V_{1}\cup V_{2}$. Since $G$ is a $2$-tree, it is chordal, and hence $\ell^{\prime}_{i}$ has a chord $e$. If $e$ is not incident with $x_{i}$, then there exists a cycle whose vertices belong to $V_{1}\cup V_{2}$, which contradicts the fact that $V_{1}\cup V_{2}$ induces a forest. Thus, $e$ is incident with $x_{i}$. Let $p$ and $q$ be vertices of $H_{i}$ such that $p$ lies inside $\ell^{\prime}_{i}$ and $q$ lies outside $\ell^{\prime}_{i}$. Then, we can find a shorter cycle than $\ell^{\prime}_{i}$ in $H_{i}$ which includes $e$ and separates $p$ and $q$. Therefore, we can eventually find a cycle of length $3$ which separates $p$ and $q$. We denote it by $\ell_{i}$.
Let $\mathcal{C}=\{\ell_{i}\mid 1\leq i\leq m\}$. Since every cycle in $\mathcal{C}$ has length $3$, any two cycles in $\mathcal{C}$ do not cross transversely on the plane, that is, the $2$-cell regions bounded by them are disjoint or one contains the other. (These cycles may share vertices and an edge.) Thus, there exists a cycle in $\mathcal{C}$ such that no other cycles in $\mathcal{C}$ embedded inside it.
Suppose that there exist two separating cycles, say $\ell_{1}$ and $\ell_{m}$, in $\mathcal{C}$ such that no other cycles in $\mathcal{C}$ are embedded inside them. If there are at least two separating $3$-cycles containing $x_{1}$ in $H_{1}$, we take $\ell_{1}$ to be the innermost one. Similarly, we take $\ell_{m}$ to be the innermost separating $3$-cycles containing $x_{m}$ in $H_{m}$. Let $y_{1}$ and $y_{m}$ be vertices inside $\ell_{1}$ and $\ell_{m}$, respectively. If $y_{1}$ belongs to $V_{2}$, we let $H=G[V_{1}\cup V_{3}\cup\{y_{1}\}]$. Since there is no vertex in $V_{3}$ inside $\ell_{1}$ and $\ell_{1}$ is the innermost separating cycle containing $x_{1}$, there is no separating cycle containing $y_{1}$ in $H$. Thus, $H$ is an outerplane graph. It follows from the inequality $|V(H)|=|V_{1}|+|V_{3}|+1\geq\lceil\frac{2n+2}{3}\rceil$ that it is a desired subgraph. Similarly, if $y_{m}$ belongs to $V_{2}$, we can find a desired subgraph. Thus, $y_{1}$ and $y_{m}$ belong to $V_{1}$. Moreover, we may assume that there is no vertex in $V_{2}$ both inside $\ell_{1}$ and inside $\ell_{m}$.
Since $G$ is a $2$-tree, every vertex of $G$ is contained in some $3$-cycle in $G$. Since every vertex inside $\ell_{1}$ belongs to $V_{1}$, they are independent, and hence they must be adjacent to the two vertices in $\ell_{1}\cap(V_{2}\cup V_{3})$. However, $\ell_{1}$ is the innermost $3$-cycle. It implies that $y_{1}$ is the only vertex inside $\ell_{1}$. Similarly, $y_{m}$ is the only vertex inside $\ell_{m}$. Thus, the induced subgraph $H^{\prime}=G[V_{2}\cup V_{3}\cup\{y_{1},y_{m}\}]$ has no separating cycle, and hence it is an outerplane graph. It follows from the inequality $|V(H^{\prime})|=|V_{2}|+|V_{3}|+2\geq\lceil\frac{2n+2}{3}\rceil$ that it is a desired subgraph.
Suppose that there is only one cycle, say $\ell_{1}$, in $\mathcal{C}$ such that no other cycles in $\mathcal{C}$ embedded inside it. Then, for any $2\leq i\leq m$, the $2$-cell region bounded by $\ell_{i}$ includes $\ell_{1}$, and hence for any two cycles in $\mathcal{C}$, one is embedded inside the other. Thus, there exists a cycle, say $\ell_{m}$, in $\mathcal{C}$ such that no other cycles in $\mathcal{C}$ embedded outside it. From the same arguments as in the case where there is no cycle in $\mathcal{C}$ embedded inside $\ell_{1}$ or $\ell_{m}$, we can obtain an induced outerplane subgraph of order at least $\lceil\frac{2n+2}{3}\rceil$: we take $\ell_{1}$ (resp. $\ell_{m}$) to be the innermost (resp. outermost) separating $3$-cycles containing $x_{1}$ (resp. $x_{m}$) in $H_{1}$ (resp. $H_{m}$). ∎
We now construct a family of outerplanar graphs which can be embedded on the plane so that this embedding achieves the lower bound given by Theorem 16. Let $G_{3}$ be the graph consisting of disjoint union of $k$ $3$-cycles $C_{1},C_{2},\ldots,C_{k}$ and three isolated vertices $v_{1},v_{2}$ and $v_{3}$. We embed $G_{3}$ on the plane so that $C_{i}$ is embedded inside $C_{i+1}$ for $11\leq i\leq k-$ 1, and $v_{2}$ lies outside $C_{k}$ and $v_{1}$ and $v_{3}$ lie inside $C_{1}$. It is easy to check that any induced outerplane subgraph of $G_{3}$ has at most $\lceil\frac{2(3k+3)+2}{3}\rceil=2k+3$ vertices. It can be easily verified that the outerplanar graphs $G_{2}=G_{3}-v_{3}$ and $G_{1}=G_{2}-v_{2}$ also achieve the lower bound given by Theorem 16.
The left and right graphs in Figure 11 represent the same outerplanar graph, which is obtained by adding some edges to $G_{3}$. The right graph shows an embedding mentioned above. (The bold lines represent edges of $k$ $3$-cycles $C_{1},C_{2},\ldots,C_{k}$)
Figure 11: A maximal outerplanar graph attaining the bound given by Theorem 16
8 Proof of Theorem 9
We construct an Eulerian plane triangulation of order $n$ with connected domination number at least $\frac{3n}{8}-1$, which shows that Conjecture M is false.
Construction: For any integer $k\geq 2$, we construct the graph $T_{k}$ as follows. We first prepare the graph $I$ shown in Figure 12, where we set $T=\{v_{1},w_{1}\}$.
$u_{1}$$u_{2}$$u_{3}$$v_{1}$$v_{2}$$v_{3}$$w_{1}$$w_{2}$$w_{3}$
Figure 12: The graph $I$, where the two vertices enclosed in circles form the set corresponding to $T_{j,i}$ defined in Remark 18.
We next prepare the graph obtained from the Cartesian product $C_{3}\Box P_{k}$ by adding diagonal edges. More precisely, we define the graph $T_{k}$ by
$V(T_{k})$ =$=\{x_{j,i},y_{j,i}\mid 1\leq j\leq k,1\leq i\leq 3\}$
$E(T_{k})$ =$=\{x_{j,i}x_{j,i+1},y_{j,i}y_{j,i+1}\mid 1\leq j\leq k,1\leq i\leq 3\}$
$\cup\{x_{j,i}y_{j,i}\mid 1\leq j\leq k,1\leq i\leq 3\}$
$\cup\{y_{j,i}x_{j+1,i}\mid 1\leq j\leq k-1,1\leq i\leq 3\}$
$\cup\{x_{j,i+1}y_{j,i}\mid 1\leq j\leq k,1\leq i\leq 3\}$
$\cup\{y_{j,i+1}x_{j+1,i}\mid 1\leq j\leq k-1,1\leq i\leq 3\},$
where we take $x_{j,4}=x_{j,1}$ and $y_{j,4}=y_{j,1}$ for each integer $1\leq j\leq k$.
$x_{1,1}$$x_{1,2}$$x_{1,3}$$x_{1,1}$$y_{1,1}$$y_{1,2}$$y_{1,3}$$y_{1,1}$$x_{2,1}$$x_{2,2}$$x_{2,3}$$x_{2,1}$$y_{2,1}$$y_{2,2}$$y_{2,3}$$y_{2,1}$$x_{k,1}$$x_{k,2}$$x_{k,3}$$x_{k,1}$$y_{k,1}$$y_{k,2}$$y_{k,3}$$y_{k,1}$
Figure 13: The graph $T_{k}$, where a copy of $I$ is embedded into each shaded triangle and the vertices enclosed in circles form (a part of) the set $W$ defined in Remark 18.
For each integer $1\leq j\leq k$ and each integer $1\leq i\leq 3$, we replace a face $F_{j,i}=x_{j,i}x_{j,i+1}y_{j,i}$ with a copy of $I$ by identifying the $3$-cycle of $F_{j,i}$ with the boundary cycle $u_{1}u_{2}u_{3}$ of $I$ so that $u_{1}=x_{j,i}$, $u_{2}=x_{j,i+1}$ and $u_{3}=y_{j,i}$. Then we denote by $I_{j,i}$ the copy of $I$ pasted in $F_{j,i}$. The resulting graph is denoted by $T_{k}$ (see Figure 13).
Let $S$ be a minimum connected dominating set of $T_{k}$. We show that $|S|\geq\frac{3n}{8}-1$. To dominate all vertices in $I$ except $u_{1},u_{2},u_{3}$, we need to choose at least two of $v_{1},v_{2},v_{3},w_{1},w_{2},w_{3}$ as a connected dominating set; otherwise, if exactly only one of them is chosen, then either at least one of them is not dominated or some $w_{i}$ chosen is not adjacent to any dominating vertex. Therefore, for each integer $1\leq j\leq k$ and each integer $1\leq i\leq 3$,
$|S\cap V(I_{j,i})\setminus\{x_{j,i},x_{j,i+1},y_{j,i}\}|\geq 2.$ (8.1)
Let $L^{x}_{j,i}=\{x_{j,i},x_{j,i+1},x_{j,i+2}\}$ and $L^{y}_{j,i}=\{y_{j,i},y_{j,i+1},y_{j,i+2}\}$. Since $G[S]$ is connected, we have one of the following for each integer $11\leq j\leq k-$ 1:
  1. 1.
    $|S\cap L^{x}_{j,i}|\geq 2$ and $|S\cap L^{y}_{j,i}|\geq 1$.
  2. 2.
    $|S\cap L^{x}_{j,i}|\geq 1$ and $|S\cap L^{y}_{j,i}|\geq 2$.
  3. 3.
    $|S\cap L^{y}_{j,i}|\geq 3$.
Otherwise, we easily see that for some $j,i$, all vertices in $S\cap I_{j,i}$ is not adjacent to any other vertex in $S\setminus I_{j,i}$ by (8.1), or $S\cap(L^{x}_{j,i}\cup L^{y}_{j,i})$ is not connected to $S\cap(L^{x}_{j+1,i}\cup L^{y}_{j+1,i})$. When $j=k$, $|S\cap(L^{x}_{k,i}\cup L^{y}_{k,i})|=2$ suffice to connect all vertices in $S\cap\bigcup_{1\leq i\leq 3}I_{k,i}$. Therefore, we can easily verify that
  • $|V(T_{k})|=24k,\ |S|=(2\cdot 3)\cdot k+3\cdot(k-1)+2=9k-1$, and then
  • $\gamma_{c}(T_{k})=|S|\geq\frac{3}{8}|V(T_{k})|-1$.
Furthermore, it can immediately be checked that $T_{k}$ is Eulerian. This completes the proof.
Remark 18
We also see that $\gamma_{c}(T_{k})=\frac{3}{8}|V(T_{k})|-1$. For each integer $1\leq j\leq k$ and each integer $1\leq i\leq 3$, let $Z_{j,i}=\{v_{j,i},w_{j,i}\}$ be the subset of $V(I_{j,i})$ that corresponds to $\{v_{1},w_{1}\}$ of $I$. Set $W=\{x_{j,i},x_{j,i+1},y_{j,i}|1\leq j\leq k-1\}\cup\{x_{k,i},x_{k,i+1}\}$. Then it is clear that $W\cup\bigcup_{1\leq j\leq k}\bigcup_{1\leq i\leq 3}Z_{j,i}$ is a connected dominating set and its cardinality is exactly $\frac{3}{8}|V(T_{k})|-1$.
Remark 19
We also see that $s_{o^{\prime}}(T_{k})=\frac{5}{8}|V(T_{k})|+1$. By Remark 18, we have $s_{o^{\prime}}(T_{k})\geq\frac{5}{8}|V(T_{k})|+1$, and then we suffice to show that $s_{o^{\prime}}(T_{k})<\frac{5}{8}|V(T_{k})|+2$. Let $O$ be the largest induced outerplane subgraph of $T_{k}$. For each $j\in\{1,2,\dots,k\}$, let $L_{j}=L^{x}_{j,i}\cup L^{y}_{j,i}$ and let $A_{j}=L_{j}\cup I_{j,i}$, where $1\leq i\leq 3$. Suppose to the contrary that $|V(O)|\geq\frac{5}{8}|V(T_{k})|+2=15k+2$. By this assumption, we have either (i) there is some $1\leq t\leq k$ such that $|V(O)\cap(A_{t})|\geq 17$ or (ii) there is some $11\leq t\leq k-$ 1 such that $|V(O)\cap(A_{t})|\geq 16$ and $|V(O)\cap(A_{\ell})|\geq 15$ for any other $\ell\neq t$. For each integer $1\leq j\leq k$ and each integer $1\leq i\leq 3$, it is clear that $|V(O)\cap I_{i,j}|\leq 4$. Therefore, in both cases (i) and (ii), we can easily find a separating cycle in $O$, which contradicts that $O$ is outerplane.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.
Acknowledgments
This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.